solve for x, y intercepts -2x +3y =8
Answer by Mathtut(3670)

3y=2x+8--->y=3/2x+8/3
:
y intercept is 8/3
:
to find x intecept set y =8/3 (0,8/3)
:
2x+8=0
:
2x=-8
:
x=-4...x intecept is -4:(-4,0)